The decomposition of XY is second order in XY and has a rate constant of 7.06×10−3 M−1⋅s−1 at a certain temperature.

PART A

If the initial concentration of XY is 0.160 M , how long will it take for the concentration to decrease to 6.80×10−2 M ?

Express your answer using two significant figures.

PART B

If the initial concentration of XY is 0.050 M, what is the concentration of XY after 45.0 s ?

Express your answer using two significant figures.

PART C

If the initial concentration of XY is 0.050 M, what is the concentration of XY after 550 s ?

Express your answer using two significant figures.
Answer #1

1/[A]    = 1/[A0] + Kt

K = 7.06×10−3 M−1⋅s−1

part-A

1/[A]    = 1/[A0] + Kt

K = 7.06×10−3 M−1⋅s−1

[A0] = 0.16M

[A]   = 6.8*10^-2 M = 0.068M

1/0.068 = 1/0.16 + 7.06*10^-3 * t

14.7   = 6.25 + 7.06*10^-3 * t

14.7-6.25    = 7.06*10^-3 * t

8.45     = 7.06*10^-3 * t

t   = 8.45/(7.06*10^-3)    = 1196.88sec>>>>answer

part-B

1/[A]    = 1/[A0] + Kt

K = 7.06×10−3 M−1⋅s−1

[A0] = 0.05M

t   = 45s

1/[A]   = 1/0.05 + 7.06*10^-3*45

1/[A]   = 20 + 0.3177

1/[A]   = 20.3177

[A]    = 1/20.3177 = 0.049M

part-C

1/[A]    = 1/[A0] + Kt

K = 7.06×10−3 M−1⋅s−1

[A0] = 0.05M

t   = 500s

1/[A]   = 1/0.05 + 7.06*10^-3*500

1/[A]   = 20 + 3.53

1/[A]   = 23.53

[A]    = 1/23.53 = 0.043M